# Open Problems, and Future Trends

### v1

@@ II-D Extended Reality
II-D1 Research Advances of XR Enabling Industrial Metaverse
@@
An intriguing concept in this realm is the utilization of an extended Metaverse agent approach based on the MiRAs cube theory[52]. This approach serves the purpose of dynamically governing virtual device control and hologram modeling, while facilitating interactions with other holograms and operators to manifest within the physical realm. Building upon this notion, Rincon et al.[53] leverage agents founded on the meta-model (MAM5) and JaCalIVE frameworks to mitigate discrepancies between simulated and real systems. Intelligent resource artifacts are incorporated, enabling seamless access to the physical world for devices. The dynamic essence of agents is further explored by Croatti et al.[54], whose architecture allows for agents to be dynamically altered at runtime, employing diverse heterogeneous agent programming techniques to modify agents within AR. In addition to system modeling, certain researchers have embraced agent-based approaches to automate XR quality testing and support experience evaluation by adopting user roles[55]. Braud et al.[56] construct an XR architecture from the operating system level, which perceives the physical world as a shared resource among applications. By implementing meticulous control mechanisms through several components such as environment understanding, specialized chip drivers, networking, interoperability, and various other interconnected elements, they achieve enhanced integration of content with the physical-digital world, thereby attaining superior performance. Drawing upon the extended Metaverse architecture, Guan et al.[51] conduct an in-depth analysis of interface design consistency within the Metaverse. They attribute inconsistencies to signal noise and propose the utilization of ”noise reduction” techniques to mitigate disparities between the Metaverse and the physical world. This approach aids in facilitating the development of interoperable interfaces. The authors further delve into the realm of seamless integration by exploring the presence of interoperable agents. They showcase the effectiveness of this approach through two early design projects, presenting interactive and tangible Metaverse applications that can be deployed across both physical and virtual realities. As interfaces of this nature continue to be researched and developed, the integration of suitable filters and controls becomes crucial in ensuring the sustained continuity of the Extended Metaverse. This integration holds the potential to serve as a valuable research direction in the long-term future.
@@
Another intriguing perspective to explore is whether the utilization of AR/VR devices can have adverse effects on workload, considering that workload predominantly reflects the input costs in production tasks. Kantowitz et al.[62] demonstrate that efficiency tends to decline as task complexity increases during HCI, even within the maximum workload capacity. To put it simply, it raises the question of whether operators need to exert additional cognitive effort to execute tasks through VR interfaces that could otherwise be performed in the real world. In this context, Xi et al.[63] conduct an investigation on the workload disparities of XR across six distinct dimensions. Their finely detailed analysis reveal that AR has a more pronounced impact on workload compared to VR, while the combination of both technologies do not escalate task difficulty. Future research endeavors can delve into exploring the mechanisms and boundary conditions of these different dimensions, enabling XR designers to develop more scientifically informed HCI systems.
II-D2 Application Stage in Industrial Production
Product Recognition and Modeling: Object detection is widely applied in XR and is an essential task for achieving the Metaverse. For example, computer vision techniques are used for real-time recognition and overlaying relevant information on the display screen, such as product specifications, usage instructions, and maintenance methods. It can also be applied to the modeling and visualization of industrial products, where real-time visual feedback helps improve the efficiency of product design, assembly, and maintenance.
Device Pose Recognition: In XR, it is often necessary to observe and recognize the actions of devices like robots in a 3D immersive factory environment and provide feedback for specific actions. For instance, on an assembly line, AR technology can be used to detect the pose of devices or components in real time, including rotation angles, positions, and sizes, thus assisting operators in correct assembly and adjustment.
II-D3 Summary
@@
TABLE III: Research on enabling technologies in the Industrial Metaverse.
@@
[19] 2022 Access management. Responding to Metaverse and decentralized Internet initiatives. Dedicated frequency bands in regional spectral bands for link transmission. Lack of simulation.
[21] 2022 Cross-enterprise collaboration. Optimizing resource sharing and utilization. A practical BC-based MEC platform is proposed for resource sharing and optimal utilization to accomplish requested offloading tasks. No detailed assessment of the proposed system is provided.
@@
[35] 2022 Base station data receiving. Ensuring the quality of VSP’s services. The authors investigate the multi-base station channel resource allocation problem in a Metaverse DT with an end-to-end cloud three-layer framework. Storage resources are not considered.
[29] 2022 Surface defect detection. DT solutions for small surface defect inspection tasks. Integration of multimodal data. Algorithms struggle to handle massive data scenarios.
[33] 2022 Base station data receiving. Reduce cloud load with lightweight DT integration with MEC. A Service Placement Algorithm Based on Merkle Tree. The baseline algorithm is not enough.
[37] 2022 Crane application. Improved efficiency of XR application development. The architecture includes a perception layer and ROS-based simulation model, and a Unity-based service layer to create XR-based digital applications. It may not be possible to apply it to other areas.
[40] 2023 Supply chain and O&M. Promoting synergy between DT and XR. The TwinXR approach is proposed, which combines information management DT and knowledge-based XR. Feature mapping is entirely dependent on shared data structure.
6G [4] 2022 Computer vision. Organic integration of edge AI and Metaverse with 6G support. Three novel edge Metaverse architectures are introduced that use 6G-enabled edge AI to address resource and computational limitations. Missing network latency and privacy issues with 6G.
@@
[65] 2022 Interaction & Assistant. Adapting to MTMM workload requirements for device heterogeneity. Various features of real-time MMMT ML workloads are discussed, and XRB ENCH and three representative executions are introduced. No consideration of cascading pipeline complexity.
[66] 2022 None. Understanding the relationship in XR’s perception of user privacy. The article analyzes the gap between user privacy perceptions and their specific behaviors across XR. Omission of audit and transparency.
@@
AI [68] 2022 Multi-Team collaboration. Lightweight Metaverse framework to support language and semantic. Summarizes common AI technologies and application development templates, as well as common functional modules and interfaces. Lack of quantitative assessment.
[69] 2022 Product design. Improve 3D modeling efficiency. AI-assisted VR modeling. Lack of baseline comparison.
[70] 2022 Product design. Memory interfaces are too restrictive in terms of bandwidth. Higher XR resolution can be achieved by stacking multiple layers of memory using 3D stacking. Lack of practical examples for deployment.
@@
[72] 2022 Product performance analysis. Improving the reliability of numerical modeling. Four ML methods are used to fabricate advanced DTs of 42SiCr alloy considering all their uncertainties and nonlinearities. The industrial scenarios involved are relatively homogenous.
II-E Artificial Intelligence
@@
II-E1 Research Advances of Artificial Intelligence Enabling Industrial Metaverse
@@
First, under AI-assisted mechanisms, XR includes technologies such as computer vision and XR computing, which can immediately capture environmental data within the field of view and convey real-time processing results to users through display devices. On one hand, research has been conducted on adaptive architectures for XR devices. Wong et al.[69] devise a system called GetWild, where the main network in the AI stage is CGAN. This stage allows users to capture images using VR devices as input to create rough 3D objects and terrain, and further perform editing operations. This technology can be effectively applied in the industrial product design process. However, this massive 3D modeling can pose resource constraints for small edge devices such as VR goggles. Even with optimizers for XR device workloads, latency issues caused by high memory requirements still exist. To address this, Yang et al.[70] employ a 3D stacking neural network accelerator architecture to evaluate energy and latency improvements. The simulator provides estimates of various hardware metrics, such as execution time (latency), energy, input/output flow, and resource utilization of prototype ML accelerators for running AR/VR models. Based on the analytical model of the simulator, operators are extracted, and expected performance indicators are calculated based on the energy table of the accelerator architecture to guide architecture design decisions for different ML models. Subsequent work can further expand the memory capacity in the X, Y, and Z directions to identify points of diminishing returns and consider the practical deployment and implementation of the platform. In addition, for VR performance estimation, Yang et al.[74] contribute a dataset (VRQ-TJU) that can be used for VR quality assessment and introduced an end-to-end 3D CNN for video quality assessment. The 3D CNN takes VR differential video patches as input and considers information between different frames. Subsequent work can analyze various features rather than just spatial distribution changes. In addition to auxiliary architecture optimization and quality assessment, AI is widely used for VR image denoising, super-resolution generation, etc., to provide operators with smaller granularity of operations and clearer interactive interfaces[75],[76].
@@
In conjunction with the integration of XR, the Metaverse architecture empowered by BC, Web 3.0, and MEC has engendered the emergence of Decentralized AI (DeAI)[77]. In contrast to Centralized AI (CeAI), which embodies a static, centralized system leading to limited flexibility and scalability, DeAI possesses the capability to decompose tasks assigned by central departments into subtasks, subsequently allocating them to the next tier of agents for processing. Through collaborative efforts, these agents ultimately converge to formulate the final solution. While intricate AI models often entail substantial computational overhead, rendering real-time execution on edge devices within the Metaverse a challenge, lightweight AI processing solutions have matured. These encompass techniques such as parameter pruning, quantization, low-precision approximation, and knowledge distillation. Lee et al.[78] have successfully applied a lightweight LSTM model in Head-mounted Displays (HMDs). The network takes a set of 10 data points sampled from motion data at intervals of 5 timestamps as input. Subsequently, online learning techniques can be applied for personalization to enhance accuracy. By virtue of lightweight edge deployment, DeAI enables the realization of intelligent device agents at the edge level. Nodes within the edge layer undertake preliminary solutions through local reasoning and subsequently coordinate and communicate with neighboring nodes, thereby facilitating resource sharing. This paradigm presently represents the optimal approach for the Industrial Metaverse.
Another facet of AI-enabling technology revolves around the advancements within AI technology itself. With the escalating popularity of generative AI models such as AI generated content (AIGC), the capacity of AI to generate content has emerged as a new frontier within the Metaverse. Consider an employee equipped with an AR device tasked with arranging equipment in a virtual environment. In this scenario, the employee may require a catalog of pre-ordered equipment, prompting the input of relevant details into an AIGC model. The model then generates equipment entities based on the given context and scenario. Notably, AIGC within the Metaverse should meet two essential criteria: first, it should support device access at any time and any location; second, it should maximize user utility within the Metaverse while satisfying user needs. Nevertheless, due to variations in task suitability among AIGCs and the non-uniform distribution of computing power across edge servers, a resource scheduling scheme that ensures reasonable allocation becomes imperative. Addressing this concern, Du et al.[79] propose an optimal decision algorithm for AIGC based on a diffusion model, used to generate ASP selection decisions. The system collects AIGC task requirements from users, where user submissions include text prompts describing the desired content. Edge servers process and schedule these requests, ensuring efficient resource utilization and optimal system performance by understanding the specific requirements of each task, including computational complexity and output quality. After collecting user requirements, the system conducts semantic analysis through a knowledge graph to identify similarities and differences between user prompts. This approach provides an effective solution for deploying AIGC services in the Metaverse. Similarly focusing on edge deployment for AIGC, Xu et al.[80] synthesize the trade-offs between reasoning latency, resource consumption, and accuracy by managing pretrained foundation models (PFMs) on edge servers using the minimal context algorithm, reducing the overall system cost. This work provides a formal mathematical expression for basic model caching and inference problems and the minimal context algorithm. Overall, the emergence of AIGC illuminates new avenues of development within the Metaverse, enhancing content diversity and alleviating repetitive production labor. However, it also gives rise to novel challenges, such as the regulation of generative content and the screening of potentially explosive content[81].
@@
AI is poised to enable intelligent networking, immersive digital worlds, inclusive user interfaces, accurate avatar creation, multilingual accessibility, and many other features within the concept of the Metaverse, which is inherently user-centric[82]. The more precise and refined the AI models are, the better the overall experience can be for users. Multilingual accessibility and automatic translation are also critical components that facilitate cross-border enterprise access to the Metaverse[83]. Nevertheless, interpretable AI is essential to enhancing brain mapping accuracy, as it enables explanations for trained models and enhances the interpretability of AI black box operations. In recent years, a variety of interpretable AI-based approaches have been introduced to improve control and monitoring of unwanted or adverse effects, such as biased decision making. Interpretable AI utilizes semantic representations to uncover relationships between different object coverages in images, thereby enhancing the interpretability of deep learning solutions. Additionally, interpretable AI can be utilized for product design and marketing, helping to analyze competitor performance and identify opportunities to outperform them[84]. By leveraging customer-generated reviews, interpretable AI can extract competitive factors for products and effectively reflect customer opinions.
II-E2 Summary

### v2

@@ II-E Extended Reality
II-E1 Research Advances of XR Enabling Industrial Metaverse
@@
An intriguing concept in this realm is the utilization of an extended Metaverse agent approach based on the MiRAs cube theory[69]. This approach serves the purpose of dynamically governing virtual device control and hologram modeling, while facilitating interactions with other holograms and operators to manifest within the physical realm. Building upon this notion, Rincon et al.[70] leverage agents founded on the meta-model (MAM5) and JaCalIVE frameworks to mitigate discrepancies between simulated and real systems. Intelligent resource artifacts are incorporated, enabling seamless access to the physical world for devices. The dynamic essence of agents is further explored by Croatti et al.[71], whose architecture allows for agents to be dynamically altered at runtime, employing diverse heterogeneous agent programming techniques to modify agents within AR. In addition to system modeling, certain researchers have embraced agent-based approaches to automate XR quality testing and support experience evaluation by adopting user roles[72]. Braud et al.[73] construct an XR architecture from the operating system level, which perceives the physical world as a shared resource among applications. By implementing meticulous control mechanisms through several components such as environment understanding, specialized chip drivers, networking, interoperability, and various other interconnected elements, they achieve enhanced integration of content with the physical-digital world, thereby attaining superior performance. Drawing upon the extended Metaverse architecture, Guan et al.[68] conduct an in-depth analysis of interface design consistency within the Metaverse. They attribute inconsistencies to signal noise and propose the utilization of “noise reduction” techniques to mitigate disparities between the Metaverse and the physical world. This approach aids in facilitating the development of interoperable interfaces. The authors further delve into the realm of seamless integration by exploring the presence of interoperable agents. They showcase the effectiveness of this approach through two early design projects, presenting interactive and tangible Metaverse applications that can be deployed across both physical and virtual realities. As interfaces of this nature continue to be researched and developed, the integration of suitable filters and controls becomes crucial in ensuring the sustained continuity of the Extended Metaverse. This integration holds the potential to serve as a valuable research direction in the long-term future.
@@
Another intriguing aspect to explore is whether the use of AR/VR devices could have adverse effects on workload, as workload primarily reflects the input costs in production tasks. Kantowitz et al. [79] demonstrate that efficiency tends to decrease as task complexity increases during HCI, even within the limits of maximum workload capacity. Simply put, this raises the question of whether operators need to exert additional cognitive effort to perform tasks through VR interfaces that could otherwise be completed in the real world. Furthermore, Zhao et al. [80] highlight that by creating multisensory experiences such as visual, auditory, and tactile sensations in XR immersive environments, learners’ emotional engagement and cognitive levels can be effectively enhanced within the Metaverse. Future research could explore the mechanisms and boundary conditions of these various dimensions, providing XR designers with a more scientifically informed approach to developing HCI systems.
II-E2 Application Stage in Industrial Production
In product recognition and modeling, object detection is widely applied in XR (Extended Reality) and is essential for realizing the Metaverse. For example, systems can leverage computer vision technologies for real-time object recognition and overlay relevant information on displays, such as product specifications, usage instructions, and maintenance methods. This technology can also be applied to industrial product modeling and visualization. Real-time visual feedback allows operators to intuitively understand and adjust products, reducing design and assembly errors and improving the efficiency of product design, assembly, and maintenance.
In device pose recognition, XR environments often require the observation and recognition of device movements, such as robots, within a 3D immersive factory setting, and the generation of feedback based on specific actions. For instance, in assembly lines, AR technology can detect the pose of devices or components in real-time, including rotation angles, positions, and dimensions, thus assisting operators in accurate assembly and adjustments.
II-E3 Summary
@@
TABLE V: Research on enabling technologies in the Industrial Metaverse.
@@
[19] 2022 Access management. Responding to Metaverse and decentralized Internet efforts. Dedicated frequency bands in regional spectral bands for link transmission. Lack of simulation.
[21] 2022 Cross-enterprise collaboration. Optimizing resource sharing and utilization. A practical BC-based MEC platform is proposed for resource sharing and optimal utilization. No detailed assessment is provided.
PPC [37] 2020 General. Reducing privacy leakage without loss of performance. DPNR with differential privacy for formal privacy guarantee. Introduce noise in complex tasks.
[35] 2023 Data sharing. Balancing privacy and utility in shared data. WGAN for privacy-preserving data sharing. Real-world deployment.
[32] 2024 Healthcare services. Ensure privacy against model poisoning attacks. A personalized FL algorithm. Scalability challenges.
[33] 2024 Cross-enterprise collaboration. Assess PFL’s vulnerability to poisoning attacks. Bayesian optimization-driven attack. Limited real-world applicability.
@@
[47] 2022 Base station data receiving. Ensuring the quality of VSP’s services. The authors study resource allocation in a Metaverse DT with a cloud framework. Storage resources are not considered.
[41] 2022 Surface defect detection. DT solutions for small surface defect inspection tasks. Integration of multimodal data. Algorithms struggle with large data.
[45] 2022 Base station data receiving. Reduce cloud load with MEC. A Service Placement Algorithm Based on Merkle Tree. Baseline algorithm is not enough.
[49] 2022 Crane application. Improved efficiency of XR application development. The architecture includes a perception layer, ROS model, and Unity service layer for XR apps. It may not apply to other areas.
[52] 2023 Supply chain and O&M. Promoting synergy between DT and XR. TwinXR combines information management DT and knowledge-based XR. Feature mapping is entirely dependent on shared data structure.
5G/6G [11] 2022 Computer vision. Organic integration of edge AI and Metaverse with 6G support. Three novel edge Metaverse architectures are introduced that use 6G-enabled edge AI to address resource and computational limitations. Missing network latency and privacy issues with 6G.
@@
[82] 2022 General. Understanding XR’s impact on user privacy. The article analyzes the gap between user privacy perceptions and their specific behaviors across XR. Omission of audit and transparency.
@@
AI [84] 2022 Product design. Boost 3D modeling efficiency. AI-assisted VR modeling. Lack of baseline.
[85] 2022 Product design. Memory interfaces have limited bandwidth. Higher XR resolution can be achieved by stacking multiple layers of memory using 3D stacking. No practical deployment examples.
@@
[87] 2022 Product analysis. Improving the reliability of numerical modeling. Four ML methods create advanced DTs for 42SiCr alloy, accounting for uncertainties and nonlinearities. The industrial scenarios are homogenous.
II-F Artificial Intelligence
@@
II-F1 Research Advances of Artificial Intelligence Enabling Industrial Metaverse
@@
First, the AI-assisted XR mechanism includes technologies such as computer vision and XR computation, enabling real-time capture of environmental data within the field of view and delivering immediate processing results to users via display devices. In terms of 3D modeling, beyond basic model generation and editing, AI technology can be utilized for automated model optimization and the generation of complex geometric structures. For example, Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) are employed to automatically generate new designs, enhancing flexibility and customization capabilities in design. On one hand, research into adapting architectures for XR devices has been conducted, with Wong et al.[84] developing a system named GetWild. The core network at the AI stage of this system is a CGAN, which allows users to capture images using VR devices as input to create rough 3D objects and terrain, followed by further editing operations. This technology holds practical applications in industrial product design. However, the extensive 3D modeling can pose resource limitations for small-scale edge devices like VR headsets. Even with workload optimizers tailored for XR devices, high memory requirements still result in latency issues. To address this, Yang et al.[85] utilize a 3D-stacked neural network accelerator architecture to evaluate energy and latency improvements. This simulator estimates various hardware metrics, such as execution time (latency), energy consumption, input/output traffic, and resource utilization of prototype ML accelerators running AR/VR models. Based on analysis models, the simulator extracts operators and calculates expected performance metrics using energy tables based on accelerator architecture, guiding architectural design decisions for different ML models. Subsequent work can further expand memory capacity along the X, Y, and Z dimensions to identify points of diminishing returns and consider the practical deployment and implementation of platforms. Regarding VR performance estimation, Yang et al.[89] also provide a dataset for VR quality assessment (VRQ-TJU) and introduced an end-to-end 3D CNN for video quality evaluation. The 3D CNN takes VR differential video patches as input and considers information across different frames. Future work can analyze multiple characteristics, not limited to spatial distribution changes. In addition to architecture optimization and quality assessment, AI is widely used for VR image denoising, super-resolution generation, and more, to offer finer operational granularity and clearer interaction interfaces, supporting the intricate needs of industrial applications[90].
@@
In addition to integration with XR, a Metaverse architecture supported by BC, Web 3.0, and MEC has further spurred the development of decentralized AI (DeAI)[91]. Centralized AI (CeAI), as a static and centralized system, suffers from low flexibility and poor scalability, making it unsuitable for cross-enterprise collaborative scenarios. On the other hand, DeAI can break down tasks issued by central departments into subtasks, which are then distributed to lower-level agents for processing, ultimately forming a collaborative final solution[92]. For example, Li et al.[93] design a verification mechanism based on smart contracts and a content caching mechanism using Stackelberg games to prevent random outcomes from AI Generated Content (AIGC) services, maximizing the benefits of Metaverse participants. While intricate AI models often entail substantial computational overhead, rendering real-time execution on edge devices within the Metaverse a challenge, lightweight AI processing solutions have matured. These encompass techniques such as parameter pruning, quantization, low-precision approximation, and knowledge distillation. Lee et al.[94] have successfully applied a lightweight LSTM model in Head-mounted Displays (HMDs). The network takes a set of 10 data points sampled from motion data at intervals of 5 timestamps as input. Subsequently, online learning techniques can be applied for personalization to enhance accuracy. By virtue of lightweight edge deployment, DeAI enables the realization of intelligent device agents at the edge level. Nodes within the edge layer undertake preliminary solutions through local reasoning and subsequently coordinate and communicate with neighboring nodes, thereby facilitating resource sharing. This paradigm presently represents the optimal approach for the Industrial Metaverse.
Another topic concerning AI enabling technologies centers around AIGC. With its advanced capabilities in digital content bundling, intelligent content editing, and smart content creation, AIGC technology is rapidly becoming the foundation for content creation and delivery, and is regarded as the “engine” driving the Metaverse[95]. Consider an employee equipped with an AR device tasked with arranging equipment in a virtual environment. In this scenario, the employee may require a catalog of pre-ordered equipment, prompting the input of relevant details into an AIGC model. The model then generates equipment entities based on the given context and scenario. Notably, AIGC within the Metaverse should meet two essential criteria: first, it should support device access at any time and any location; second, it should maximize user utility within the Metaverse while satisfying user needs. Nevertheless, due to variations in task suitability among AIGCs and the non-uniform distribution of computing power across edge servers, a resource scheduling scheme that ensures reasonable allocation becomes imperative. Addressing this concern, Zeng et al.[96] consider multimodal parallel offloading scenarios, minimizing latency in wireless environments during multimodal AIGC service parallel offloading by using multimodal data RP segmentation and parallel load migration. Lin et al.[97] explore the joint deployment architecture of SemCom and AIGC, including a stepwise workflow to capture integration and coordination benefits, concluding with a discussion of several potential use cases. In summary, the emergence of AIGC has brought new promise to the development of the Metaverse, such as diversifying content and reducing repetitive production labor. Additionally, Chen et al.[98] highlight that AIGC can complement other enabling technologies. For instance, integrating AIGC into DT networks can help explore network behaviors such as routing, resource allocation, and network fault diagnosis[99]. Conversely, DT can enhance AIGC performance by optimizing data collection strategies, such as node deployment strategies, sampling frequency, and incentive mechanisms, thus improving data collection efficiency and alleviating bandwidth pressure through physical environment simulations. Of course, future work should not only continue to develop the advantages and potential of AIGC but also pay attention to regulating generated content and filtering explosive content[100].
@@
AI is poised to enable intelligent networking, immersive digital worlds, inclusive user interfaces, accurate avatar creation, multilingual accessibility, and many other features within the concept of the Metaverse, which is inherently user-centric[101]. The more precise and refined the AI models are, the better the overall experience can be for users. Multilingual accessibility and automatic translation are also critical components that facilitate cross-border enterprise access to the Metaverse. Nevertheless, interpretable AI is essential to enhancing brain mapping accuracy, as it enables explanations for trained models and enhances the interpretability of AI black box operations. In recent years, a variety of interpretable AI-based approaches have been introduced to improve control and monitoring of unwanted or adverse effects, such as biased decision making. Interpretable AI utilizes semantic representations to uncover relationships between different object coverages in images, thereby enhancing the interpretability of deep learning solutions. Additionally, interpretable AI can be utilized for product design and marketing, helping to analyze competitor performance and identify opportunities to outperform them[102]. By leveraging customer-generated reviews, interpretable AI can extract competitive factors for products and effectively reflect customer opinions.
II-F2 Summary